Arizona is known to have extreme temperatures during its long summers. Jose purchased a car in Flagstaff and decided to drive down to phoenix to show off his new car to his parents. Before leaving the cool 80-degree weather, and driving down to the 110-degree Phoenix city, he stopped to add air to his tires. He set all of his tires to 41psi and made his way down to phoenix. Before driving off, he received a text message from his mom, to make sure he checks his oil level, clean his mirrors and reduce his tire pressure. Why would his mom advise him to reduce the tire pressure? Your explanation should include before and after diagrams of the tire and gas particles inside, and evidence of support from the figures, principles of gas laws and thermodynamics.
1
Expert's answer
2020-09-14T08:20:02-0400

The tire pressure will increase in Phoenix city due to temperature change, so Jose’s mom advise him to reduce the tire pressure.

P1 = 41 psi

P2 = unknown

V1 = V2 = constant tire volume

T1 = 299.8 K

T2 = 316.4 K

Gay-Lussac’s law:

"\frac{P_1}{T_1} = \frac{P_2}{T_2}"

"P_2 = \frac{P_1\times T_2}{T_1}"

"P_2 = \frac{41\times316.4}{299.8} = 43\; psi"

Increasing temperature increases collisions and increases pressure.



According to the kinetic theory, the pressure of a gas results from the bombardment

of container walls by molecules. Both the concentration of molecules (number per

unit volume) and the average speed of the molecules are factors in determining this pressure. Molecular concentration and average speed determine the frequency of collisions with the wall. Average molecular speed determines the average force of a collision. If you raise the temperature, you increase the average molecular speed. The average force of a collision increases. If all other factors remained fixed, the pressure would increase.